A bag contains dimes and nickels. There are 18 coins in the bag. The value of the coins is $1.25. How many nickels are in the bag?
Answer by richwmiller(17219)

x+y=18,
5*x+10*y=125
x=18-y
5*(18-y)+10*y=125
90-5y+10*y=125
5*y=35
y=7
x=18-y
x=11 at 5 and y=7 at 10
check
11+7=18
5*x+10*y=125
5*11+10*7=125
55+70=125
125=125
ok